OpenStudy (anonymous):

x+3y=1 -5x+4y=-15 linear system by using elimination

OpenStudy (bahrom7893):

eliminate x from the secon equation by changing the first equation to: x = 1 - 3y

OpenStudy (anonymous):

x + 3y = 1 --->(5)x + 3y = 1 -5x + 4y =- 15 ----------------- 5x + 15y = 5 (result of multiplying by 5) -5x + 4y = - 15 -----------------add 0 + 19y = - 10 19y = -10 y = - 10/19 now sub - 10/19 in for y x + 3y = 1 x + 3(-10/19) = 1 x - 30/19 = 1 x = 30/19 + 1 x = 30/19 + 19/19 x = 49/19 check... x + 3y = 1 49/19 + 3(-10/19) = 1 49/19 - 30/19 = 1 19/19 = 1 1 = 1 (correct) x = 49/19 and y = -10/19

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

not a problem...glad i could help :)